OpenStudy (anonymous):

how do i graph the function given below? y=x^2-6x-7

OpenStudy (radar):

Make a table and assign some values to x, make them simple both negative and positive and obtain a corresponsing y value.

OpenStudy (anonymous):

but i dont know how to graph it

OpenStudy (anonymous):

Make a x|y chart. Have two columns, one of which being x and the other y. Then choose some points, like 0 for example to create (x,y) coordinate. Then plug the 0 into everywhere there is an X. So its gonna look like this: y = 0^2-6(0)-7 y = 0-0-7 y=0-7 y=-7 So one of your points is (0, -7). So you don't move on the x axis, but you do move 7 lines down the y axis. Just do the same steps with other values for x to get your values for y.

OpenStudy (anonymous):